# Open Vision Reasoner: Transferring Linguistic Cognitive Behavior for Visual Reasoning

### v1

@@ 1 Introduction
To this end, we introduce a robust two-stage methodology designed to first instill linguistic cognitive patterns and then activate them for visual reasoning. Our process begins with a large-scale cold start, fine-tuning Qwen2.5-VL-7B on over 2 million examples to build a strong foundation. This is followed by a prolonged reinforcement learning phase under the Open-Reasoner-Zero [24] framework, leveraging over 0.3 million mixed-modality examples. To the best of our knowledge, this represents the largest open-source RL practice on this model. The resulting model, Open-Vision-Reasoner (OVR), validates our approach by achieving strong performance across both language and multimodal benchmarks. As shown in Fig. 1, it achieves 63.5% on AIME2024 and 95.3% on MATH500 for math reasoning, as well as 51.8% on MathVision and 54.6% on MathVerse for visual reasoning.
To further trace the transfer and evolution of cognitive patterns throughout training, we develop a in-depth visual cognitive behavior analysis. Three central insights are worth highlighting: (1) Behavior transfer emerges remarkably early in cold start, driven by linguistic patterns encoding mental imagery [25, 26] as illustrated in Fig. 11. (2) Cold start broadly memorizes diverse visual cognitive behaviors, while RL critically discerns and scales up effective patterns. (3) Transfer follows a strategic path, favoring behaviors with high utility such as visual reflection. These findings deepen the understanding on visual intelligence scaffolded by linguistic reasoning [27].
@@ 2 Cognitive Behavior Preliminaries
Based on this, we investigate the transfer of cognitive behaviors from language to vision. We define the visual extensions of the aforementioned behaviors—visual reflection, divide-and-conquer, visual verification, and goal-driven visual tracing. Their formal definitions, examples, and corresponding linguistic counterparts are provided in Table 1, while Fig. 2 presents a multimodal example encompassing both linguistic and visual cognitive behaviors. In the following sections, we present a simple yet effective MLLM training pipeline comprising a linguistic cold start followed by multimodal RL (Section 3), and systematically analyze the transfer and scaling of these visual cognitive behaviors (Section 5.2).
@@ 3 Open-Vision-Reasoner
In this section, we introduce Open-Vision-Reasoner (OVR), a strong multimodal reasoning model build from Qwen2.5-VL-7B [28], from perspectives of training pipeline (Section 3.1), RL algorithm (Section 3.2) and data construction (Section 3.3).
@@ 4 Experiments
In this section, we first elaborate our implementation of Open-Vision-Reasoner (OVR). Then, we present superior performance across textual benchmarks (Section 4.2) and multimodal scenerios (Section 4.3).
@@ 4.3 Superior Visual Reasoning Abilities
As shown in Table 3, our model sets a new breakthrough for 7B models in visual reasoning. It is the first post-trained Qwen2.5-VL-7B-based model to surpass the 50% performance on MathVision, while also achieving state-of-the-art results among 7B models on DynaMath and MathVerse. This strong overall performance is further underscored by a substantial gain on MMMU-Pro (+7.2% over prior SOTA methods). These results indicate that reasoning capabilities acquired through language training can effectively transfer to multimodal tasks, leading to notable improvements in visual reasoning.
@@ 5.1 Analysis of Training Dynamics
In this section, we present a comprehensive overview of the training dynamics as illustrated in Fig. 3, and provide a detailed analysis of how text and multi-modal reasoning metrics evolve throughout the process as shown in Fig. 4.
During the initial cold-start phase (Fig. 3 (a)), the model’s loss rapidly descends to below 0.5. Subsequently, across multiple training epochs, the loss exhibits a step-wise, gradual decrease. In parallel, we observe a corresponding surge in performance across all benchmarks (Fig. 4), which first ascend sharply before transitioning to a phase of slower, more incremental improvement toward their peak. A noteworthy observation is that the aggressive training strategy detailed in Section 4.1—employing a large batch size in concert with a high learning rate—proves to be essential. This approach is critical for breaking the model’s inherent constraints, thereby successfully imbuing it with new cognitive paradigms and sculpting a more favorable landscape for reinforcement learning. It is a prerequisite that enables our model, which originates from an instruction-tuned base, to ultimately achieve text performance that is comparable to, or even surpasses models initialized from base [28] or math-specific checkpoints [78].
Furthermore, Fig. 3 (b) reveals how the model’s reward and average token length in the RL phase steadily advance from an initial 7k to exceed 12k. Owing to the stability of the training configuration inherited from Open-Reasoner-Zero [24], OVR is successfully trained on a diverse corpus of over 20 multi-modal and language-only datasets without encountering any training collapse or performance degradation. Critically, whenever the token length begins to plateau or even decline, we strategically switch to a longer context length, which invariably catalyzes the next wave of rapid reward growth. Fig. 4 captures the coincident yet unsurprising convergent growth trajectory shared by all eight reasoning benchmarks, spanning both text and multi-modal domains, as they progressively ascend towards their zenith amidst fluctuations.
@@ 5.2 Multimodal Cognitive Behavior Analysis
Recent studies have highlighted the emergence of cognitive behaviors in MLLMs during visual reasoning tasks—phenomena often dubbed “visual aha moments” [79, 21]. In this work, we move beyond plain observations and systematically investigate how these behaviors are transferred from their linguistic counterparts. Our analysis centers on the four pivotal visual cognitive behaviors introduced in Section 2 which are drawn from foundational research on cognitive patterns [13]. To quantify this process, we employ GPT-4o [60] to analyze the emergence of each behavior within the inference traces of our OVR model.
@@
Following Section 5.1, we tracked the dynamics of visual reflection, a significant behavior mentioned in previous studys [80, 81], throughout OVR’s training. As depicted in Fig. 5, this vision-specific behavior emerges in significant quantities from the very beginning of the cold-start phase and fluctuates throughout subsequent training steps. Strikingly, we observed that even in linguistic problems, DeepSeek-R1’s responses [7] frequently exhibited signs of mental imagery [25, 26] as shown in Fig. 11(a). The model appeared to construct internal visualizations to support mathematical reasoning, often articulated through phrases such as “let me visualize…” or “let me see the image.” Once this linguistic scaffolding was introduced into our MLLM, these mental images were rapidly grounded in actual visual input, enabling their rapid and effective generalization within OVR.
@@
We further investigate how cognitive behaviors scale during large-scale RL. As shown in Fig. 5(a), after an initial, rapid instillation of patterns during the aggressive cold-start phase, their prevalence is first suppressed then amplified to unprecedented levels during multimodal RL. This counter-intuitive dynamic suggests a clear division of labor: the cold-start phase learns broadly, indiscriminately memorizing all available patterns. In contrast, RL discerns critically, acting as a strategic filter for the crucial tokens [82] and scaling up pivotal behaviors. This process of RL—discarding the dross to select the essence—is significant for achieving superior generalization.
@@
To analyze the transition from linguistic to visual cognition, we track the emergence and transfer rates (detailed in Section C.1) of four core cognitive behaviors across both language and vision modalities. As shown in Fig. 5(b), the emergence of backtracking and verification steadily increases across training stages, underscoring their growing importance. Among these, the transfer rate of backtracking shows consistent growth—from 2.5% to 17.3%—while verification exhibits near-zero transfer throughout both the cold-start and RL phases. This indicates that transfer is a strategic process, for which we posit two potential explanations: (1) Backtracking transfers more readily due to DeepSeek-R1’s [7] inherent “mental imagination” capabilities, while verification, lacking a direct linguistic precursor, is more difficult for the MLLM to internalize. (2) Mirroring how humans naturally and instinctively process visual information [80], backtracking is a more fundamental component of complex visual reasoning, making its amplification a higher priority during the strategic RL phase. We will investigate these hypotheses in greater depth in our future work.
@@ 5.3 Beyond Behavior: Visual Perception Analysis and Future Work
We evaluated both stages of OVR, along with the base model Qwen2.5-VL-7B, on a comprehensive set of multimodal benchmarks targeting visual perception and recognition (MMBench [83], BLINK [84], MMStar [85], HallusionBench [86], POPE [87], RealWorldQA [88], PhyX [89], MME [90], MMVet [91]). As shown in Table 4, performance steadily improves across tasks such as MMBench and PhyX, underscoring the effectiveness of our training paradigm. The cold-start model shows declines on several tasks, notably increased hallucinations [92, 93], likely due to token distribution shifts from large-scale linguistic data [87]. However, the regained performance on benchmarks such as MMBench and BLINK demonstrate that long-term multimodal RL can effectively mitigate these issues by discerning perceptual capabilities that are critically for multimodal tasks. Looking ahead, degradation from cold start can be mitigated either by incorporating the linguistic data into model pretraining [94, 95], or by introducing more multimodal supervision during the cold start to establish a stronger visual foundation.
@@
Throughout the multimodal RL, we observed a strong correlation between the reward and the average response length in Fig. 3, which is a finding consistent with prior practices [7, 24]. This reinforces response length as an effective reward proxy, indicative of a scaling property tied to reasoning depth and computational resources. However, when focusing on specific discriminative perceptual tasks like OCR and counting, we observe a clear divergence. As shown in Fig. 6, while the reward can be effectively increased, the average response length remains largely stagnant.
@@
Model MMBench BLINK MMStar HallusionBench POPE RealWorldQA PhyX MME MMVet
@@
Qwen2.5-VL $85.3$ $84.6$ $53.7$ $62.1$ $49.0$ $86.1$ $69.3$ $37.5$ $1659.7$ $63.9$
   +Cold Start $86.1$ $82.1$ $51.5$ $62.4$ $55.0$ $82.5$ $63.1$ $47.7$ $1549.8$ $61.8$
    +Multimodal RL $86.6$ $84.2$ $54.1$ $62.7$ $53.6$ $83.2$ $65.5$ $50.0$ $1559.1$ $63.6$
@@ Appendix A Cold-Start Data Curation Details
As mentioned in Section 3.3, a critical component of our initial policy development is the curation of the cold-start SFT dataset. This stage serves as the foundation for subsequent learning, particularly in shaping the model’s ability to exhibit structured reasoning and cognitive behavior. To this end, we adopt a multi-stage curation pipeline consisting of data collection, filtering, cleaning, and strategic reweighting.
@@
We begin by assembling a broad set of prompt-response pairs that span diverse reasoning domains. These include math, science, and general logical reasoning tasks such as puzzles, deductive tasks, and constraint satisfaction problems. Our sources include a mix of public datasets as illustrated in Section 3.3.
@@ Appendix C Details for Coginitive Behavior Evaluation
In this section, we detail definitions of metrics in cognitive behavior analysis (Section 5.2) and show the prompt for evaluation.
@@ C.1 Behavior Transfer Rate
To quantify how well language-acquired cognitive behaviors generalize to the visual modality, we define the Behavior Transfer Rate (BTR) for each behavior type introduced in Fig. 5. BTR is calculated as the ratio between the emergence rate of visual behaviors and that of their linguistic counterparts. Formally, we compute the Cognitive Behavior Emergence Rate in the visual modality ($\text{CBR}_{\text{v}}$) and in the language modality ($\text{CBR}_{\text{l}}$), both evaluated on the multimodal benchmark MathVision (mini). The BTR is then defined as:
@@ C.2 Evaluation Prompt
We design prompts for the LLM-based evaluation.  Fig. 7 showcases the prompt template for the cognitive behavior Backtracking as an example.
@@ Appendix D More Case Studies
We present examples of OVR’s reasoning process during complex multimodal tasks. In Fig. 10, the model begins by perceiving the image and applying visual divide-and-conquer to interpret each emoji individually. It then engages in subgoal setting, breaking down the task to explore multiple candidate answers based on each emoji. Throughout the process, the model exhibits backtracking, revising earlier hypotheses by cross-verifying them with elements in the image. The reasoning process concludes with a final, well-justified prediction.
@@
Figure 11: Case Study on DeepSeek-R1. Red highlights the mental imagery mentioned in Section 5.2.

### v2

@@ 1 Introduction
To this end, we introduce a robust two-stage methodology designed to first instill linguistic cognitive patterns and then activate them for visual reasoning. Our process begins with a large-scale cold start, fine-tuning Qwen2.5-VL-7B on over 2 million examples to build a strong foundation. This is followed by a prolonged reinforcement learning phase under the Open-Reasoner-Zero [24] framework, leveraging over 0.3 million mixed-modality examples. To the best of our knowledge, this represents the largest open-source RL practice on this model. The resulting model, Open-Vision-Reasoner (OVR), validates our approach by achieving strong performance across both language and multimodal benchmarks. As shown in Fig.˜1, it achieves 63.5% on AIME2024 and 95.3% on MATH500 for math reasoning, as well as 51.8% on MathVision and 54.6% on MathVerse for visual reasoning.
To further trace the transfer and evolution of cognitive patterns throughout training, we develop a in-depth visual cognitive behavior analysis. Three central insights are worth highlighting: (1) Behavior transfer emerges remarkably early in cold start, driven by linguistic patterns encoding mental imagery [25, 26] as illustrated in Fig.˜11. (2) Cold start broadly memorizes diverse visual cognitive behaviors, while RL critically discerns and scales up effective patterns. (3) Transfer follows a strategic path, favoring behaviors with high utility such as visual reflection. These findings deepen the understanding on visual intelligence scaffolded by linguistic reasoning [27].
@@ 2 Cognitive Behavior Preliminaries
Based on this, we investigate the transfer of cognitive behaviors from language to vision. We define the visual extensions of the aforementioned behaviors—visual reflection, divide-and-conquer, visual verification, and goal-driven visual tracing. Their formal definitions, examples, and corresponding linguistic counterparts are provided in Table˜1, while Fig.˜2 presents a multimodal example encompassing both linguistic and visual cognitive behaviors. In the following sections, we present a simple yet effective MLLM training pipeline comprising a linguistic cold start followed by multimodal RL (Section˜3), and systematically analyze the transfer and scaling of these visual cognitive behaviors (Section˜5.2).
@@ 3 Open-Vision-Reasoner
In this section, we introduce Open-Vision-Reasoner (OVR), a strong multimodal reasoning model build from Qwen2.5-VL-7B [28], from perspectives of training pipeline (Section˜3.1), RL algorithm (Section˜3.2) and data construction (Section˜3.3).
@@ 4 Experiments
In this section, we first elaborate our implementation of Open-Vision-Reasoner (OVR). Then, we present superior performance across textual benchmarks (Section˜4.2) and multimodal scenerios (Section˜4.3).
@@ 4.3 Superior Visual Reasoning Abilities
As shown in Table˜3, our model sets a new breakthrough for 7B models in visual reasoning. It is the first post-trained Qwen2.5-VL-7B-based model to surpass the 50% performance on MathVision, while also achieving state-of-the-art results among 7B models on DynaMath and MathVerse. This strong overall performance is further underscored by a substantial gain on MMMU-Pro (+7.2% over prior SOTA methods). These results indicate that reasoning capabilities acquired through language training can effectively transfer to multimodal tasks, leading to notable improvements in visual reasoning.
@@ 5.1 Analysis of Training Dynamics
In this section, we present a comprehensive overview of the training dynamics as illustrated in Fig.˜3, and provide a detailed analysis of how text and multi-modal reasoning metrics evolve throughout the process as shown in Fig.˜4.
During the initial cold-start phase (Fig.˜3 (a)), the model’s loss rapidly descends to below 0.5. Subsequently, across multiple training epochs, the loss exhibits a step-wise, gradual decrease. In parallel, we observe a corresponding surge in performance across all benchmarks (Fig.˜4), which first ascend sharply before transitioning to a phase of slower, more incremental improvement toward their peak. A noteworthy observation is that the aggressive training strategy detailed in Section˜4.1—employing a large batch size in concert with a high learning rate—proves to be essential. This approach is critical for breaking the model’s inherent constraints, thereby successfully imbuing it with new cognitive paradigms and sculpting a more favorable landscape for reinforcement learning. It is a prerequisite that enables our model, which originates from an instruction-tuned base, to ultimately achieve text performance that is comparable to, or even surpasses models initialized from base [28] or math-specific checkpoints [78].
Furthermore, Fig.˜3 (b) reveals how the model’s reward and average token length in the RL phase steadily advance from an initial 7k to exceed 12k. Owing to the stability of the training configuration inherited from Open-Reasoner-Zero [24], OVR is successfully trained on a diverse corpus of over 20 multi-modal and language-only datasets without encountering any training collapse or performance degradation. Critically, whenever the token length begins to plateau or even decline, we strategically switch to a longer context length, which invariably catalyzes the next wave of rapid reward growth. Fig.˜4 captures the coincident yet unsurprising convergent growth trajectory shared by all eight reasoning benchmarks, spanning both text and multi-modal domains, as they progressively ascend towards their zenith amidst fluctuations.
@@ 5.2 Multimodal Cognitive Behavior Analysis
Recent studies have highlighted the emergence of cognitive behaviors in MLLMs during visual reasoning tasks—phenomena often dubbed “visual aha moments” [79, 21]. In this work, we move beyond plain observations and systematically investigate how these behaviors are transferred from their linguistic counterparts. Our analysis centers on the four pivotal visual cognitive behaviors introduced in Section˜2 which are drawn from foundational research on cognitive patterns [13]. To quantify this process, we employ GPT-4o [60] to analyze the emergence of each behavior within the inference traces of our OVR model.
@@
Following Section˜5.1, we tracked the dynamics of visual reflection, a significant behavior mentioned in previous studys [80, 81], throughout OVR’s training. As depicted in Fig.˜5, this vision-specific behavior emerges in significant quantities from the very beginning of the cold-start phase and fluctuates throughout subsequent training steps. Strikingly, we observed that even in linguistic problems, DeepSeek-R1’s responses [7] frequently exhibited signs of mental imagery [25, 26] as shown in Fig.˜11(a). The model appeared to construct internal visualizations to support mathematical reasoning, often articulated through phrases such as “let me visualize…” or “let me see the image.” Once this linguistic scaffolding was introduced into our MLLM, these mental images were rapidly grounded in actual visual input, enabling their rapid and effective generalization within OVR.
@@
We further investigate how cognitive behaviors scale during large-scale RL. As shown in Fig.˜5(a), after an initial, rapid instillation of patterns during the aggressive cold-start phase, their prevalence is first suppressed then amplified to unprecedented levels during multimodal RL. This counter-intuitive dynamic suggests a clear division of labor: the cold-start phase learns broadly, indiscriminately memorizing all available patterns. In contrast, RL discerns critically, acting as a strategic filter for the crucial tokens [82] and scaling up pivotal behaviors. This process of RL—discarding the dross to select the essence—is significant for achieving superior generalization.
@@
To analyze the transition from linguistic to visual cognition, we track the emergence and transfer rates (detailed in Section˜C.1) of four core cognitive behaviors across both language and vision modalities. As shown in Fig.˜5(b), the emergence of backtracking and verification steadily increases across training stages, underscoring their growing importance. Among these, the transfer rate of backtracking shows consistent growth—from 2.5% to 17.3%—while verification exhibits near-zero transfer throughout both the cold-start and RL phases. This indicates that transfer is a strategic process, for which we posit two potential explanations: (1) Backtracking transfers more readily due to DeepSeek-R1’s [7] inherent “mental imagination” capabilities, while verification, lacking a direct linguistic precursor, is more difficult for the MLLM to internalize. (2) Mirroring how humans naturally and instinctively process visual information [80], backtracking is a more fundamental component of complex visual reasoning, making its amplification a higher priority during the strategic RL phase. We will investigate these hypotheses in greater depth in our future work.
@@ 5.3 Beyond Behavior: Visual Perception Analysis and Future Work
We evaluated both stages of OVR, along with the base model Qwen2.5-VL-7B, on a comprehensive set of multimodal benchmarks targeting visual perception and recognition (MMBench [83], BLINK [84], MMStar [85], HallusionBench [86], POPE [87], RealWorldQA [88], MME [89], MMVet [90]). As shown in Table˜4, performance steadily improves across tasks such as MMBench, underscoring the effectiveness of our training paradigm. The cold-start model shows declines on several tasks, notably increased hallucinations [91, 92], likely due to token distribution shifts from large-scale linguistic data [87]. However, the regained performance on benchmarks such as MMBench and BLINK demonstrate that long-term multimodal RL can effectively mitigate these issues by discerning perceptual capabilities that are critically for multimodal tasks. Looking ahead, degradation from cold start can be mitigated either by incorporating the linguistic data into model pretraining [93, 94], or by introducing more multimodal supervision during the cold start to establish a stronger visual foundation.
@@
Throughout the multimodal RL, we observed a strong correlation between the reward and the average response length in Fig.˜3, which is a finding consistent with prior practices [7, 24]. This reinforces response length as an effective reward proxy, indicative of a scaling property tied to reasoning depth and computational resources. However, when focusing on specific discriminative perceptual tasks like OCR and counting, we observe a clear divergence. As shown in Fig.˜6, while the reward can be effectively increased, the average response length remains largely stagnant.
@@
Model MMBench BLINK MMStar HallusionBench POPE RealWorldQA MME MMVet
@@
Qwen2.5-VL $85.3$ $84.6$ $53.7$ $62.1$ $49.0$ $86.1$ $69.3$ $1659.7$ $63.9$
     +Cold Start $86.1$ $82.1$ $51.5$ $62.4$ $55.0$ $82.5$ $63.1$ $1549.8$ $61.8$
        +Multimodal RL $86.6$ $84.2$ $54.1$ $62.7$ $53.6$ $83.2$ $65.5$ $1559.1$ $63.6$
@@ Appendix A Cold-Start Data Curation Details
As mentioned in Section˜3.3, a critical component of our initial policy development is the curation of the cold-start SFT dataset. This stage serves as the foundation for subsequent learning, particularly in shaping the model’s ability to exhibit structured reasoning and cognitive behavior. To this end, we adopt a multi-stage curation pipeline consisting of data collection, filtering, cleaning, and strategic reweighting.
@@
We begin by assembling a broad set of prompt-response pairs that span diverse reasoning domains. These include math, science, and general logical reasoning tasks such as puzzles, deductive tasks, and constraint satisfaction problems. Our sources include a mix of public datasets as illustrated in Section˜3.3.
@@ Appendix C Details for Coginitive Behavior Evaluation
In this section, we detail definitions of metrics in cognitive behavior analysis (Section˜5.2) and show the prompt for evaluation.
@@ C.1 Behavior Transfer Rate
To quantify how well language-acquired cognitive behaviors generalize to the visual modality, we define the Behavior Transfer Rate (BTR) for each behavior type introduced in Fig.˜5. BTR is calculated as the ratio between the emergence rate of visual behaviors and that of their linguistic counterparts. Formally, we compute the Cognitive Behavior Emergence Rate in the visual modality ($\text{CBR}_{\text{v}}$) and in the language modality ($\text{CBR}_{\text{l}}$), both evaluated on the multimodal benchmark MathVision (mini). The BTR is then defined as:
@@ C.2 Evaluation Prompt
We design prompts for the LLM-based evaluation. Fig.˜7 showcases the prompt template for the cognitive behavior Backtracking as an example.
@@ Appendix D More Case Studies
We present examples of OVR’s reasoning process during complex multimodal tasks. In Fig.˜10, the model begins by perceiving the image and applying visual divide-and-conquer to interpret each emoji individually. It then engages in subgoal setting, breaking down the task to explore multiple candidate answers based on each emoji. Throughout the process, the model exhibits backtracking, revising earlier hypotheses by cross-verifying them with elements in the image. The reasoning process concludes with a final, well-justified prediction.
@@
Figure 11: Case Study on DeepSeek-R1. Red highlights the mental imagery mentioned in Section˜5.2.